### v1

@@ Abstract.
The application of Vision-Language Models (VLMs) in remote sensing (RS) image understanding has achieved notable progress, demonstrating the basic ability to recognize and describe geographical entities. However, existing RS-VLMs are mostly limited to image-level and region-level tasks, lacking the capability to handle pixel-level tasks and performing poorly in small-object recognition scenarios. Moreover, RS-VLMs consume significant computational resources when processing high-resolution RS images, further restricting their practical applicability. In this context, we propose GeoMag (Geographical Magnifier), an end-to-end general-purpose large model framework for RS. GeoMag dynamically focuses the attention scope based on prompt semantics to effectively perform remote sensing image parsing across multiple levels of granularity. This method introduces Task-driven Multi-granularity Resolution Adjustment (TMRA) and Prompt-guided Semantic-aware Cropping (PSC), which adaptively reduce the spatial resolution of task-irrelevant regions while enhancing the visual representation of task-relevant areas. This approach improves the model’s perception of critical target regions, suppresses background redundancy, and reduces the computational cost of interpreting high-resolution RS imagery. Extensive comparative experiments on 10 benchmarks demonstrate that GeoMag not only excels in handling pixel-level tasks but also maintains competitive performance across tasks of other granularities compared to existing RS-VLMs.
VLMs, Remote Sensing Image Parsing, Pixel-level Tasks, Attention Mechanism.
*Corresponding author
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; 2025; GeoMagisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Image searchccs: Information systems Multimedia and multimodal retrieval
@@ 1. Introduction
Figure 1. Overview of GeoMag’s capabilities. In addition to understanding both global and local image information for performing image-level tasks (green box) and region-level tasks (blue box), GeoMag extends the functionality of existing RS-VLMs by enabling pixel-level task processing—an ability typically absent in current models. As illustrated by the red box, GeoMag can automatically generate detailed descriptions of the image and produce segmentation masks for key geographical entities mentioned in the text. It can also perform user-directed segmentation based on specified prompts.
Remote sensing (RS) imagery, as a vital source of Earth observation data, serves scientific research, decision-making support, and the monitoring and evaluation of Sustainable Development Goals (SDGs) (Li et al., 2023). Therefore, effectively understanding the rich semantic content embedded in RS images becomes a fundamental step toward unlocking the value of RS data and promoting the intelligent development of earth science research.
@@ 2. Related Works
Pixel-level Parsing. In computer vision, pixel-level RS image parsing tasks mainly include Referring Segmentation and Open Vocabulary Segmentation (Zhou et al., 2024; Pan et al., 2024; Liu et al., 2024c). To enable large language models (LLMs) to perform pixel-level tasks, PixelLM (Ren et al., 2024) and LISA (Lai et al., 2024) introduced a [SEG] token into the vocabulary of the LLM as a conditional input, allowing the decoder to generate segmentation masks from textual descriptions. GSVA (Xia et al., 2024) further introduced a [REJ] token to explicitly exclude irrelevant or unwanted targets, enabling the model to follow user intent more precisely. GLaMM (Rasheed et al., 2024) proposed a novel pixel-level task called Grounded Conversation Generation (GCG), aiming to identify and localize specific regions or objects in images through human-machine dialogue, thereby opening a promising new direction for pixel-level task research. However, these studies were conducted on natural images, which differ significantly from RS images in terms of semantic content and visual characteristics. Therefore, directly transferring such methods often fails to yield satisfactory results (Wang et al., 2023), and specialized optimization and adaptation for RS imagery are necessary.
@@ 3.2. MagCrop
To more effectively address challenges in remote sensing image parsing—such as varying granularity, sparse targets, and uneven spatial distributions—this work proposes an image preprocessing strategy named MagCrop. MagCrop reduces resource consumption (e.g., memory, GPU usage, and time) and enhances model response accuracy through the following principles: (1) For tasks emphasizing global features, it appropriately reduces image resolution. (2) For tasks with explicit spatial references (e.g., specific locations or objects), it preserves the resolution of task-relevant regions while compressing irrelevant areas. MagCrop integrates two key mechanisms: Task-driven Multi-granularity Resolution Adjustment (TMRA) and Prompt-guided Semantic-aware Cropping (PSC). These mechanisms help maintain the resolution of task-relevant regions while effectively reducing computational redundancy and emphasizing critical semantic content.
@@ 3.2.1. Task-driven Multi-granularity Resolution Adjustment
To address these issues, this study proposes a Task-driven Multi-granularity Resolution Adjustment (TMRA) strategy. This strategy dynamically adjusts the input image resolution based on the characteristics of each task, ensuring a balance between accuracy and computational efficiency.
@@
By adaptively adjusting the image input, TMRA significantly improves GeoMag’s efficiency across tasks of different granularities, reduces the demand for computational resources.
@@ 3.2.2. Prompt-guided Semantic-aware Cropping
Current VLMs exhibit limited capability in perceiving and accurately parsing small visual details (Zhang et al., 2025; Kuckreja et al., 2024). To address the challenges RS-VLMs face in detecting small objects and to enhance their ability to focus on specific regions, GeoMag introduces Prompt-guided Semantic-aware Cropping (PSC). This strategy employs a gradient-weighted attention map generation method to query key regions in the image, emphasizing the deep semantic correlation between the prompt and the image content.
@@
By combining gradient weighting with the attention map generation strategy, PSC enhances RS-VLM’s ability to focus on key regions and small objects within the image.
@@ 5. Conclusion
This study proposes GeoMag, a RS-VLM capable of handling multi-granularity RS image parsing tasks. By introducing components such as a Pixel Decoder, GeoMag extends its capabilities to pixel-level processing. It further presents MagCrop, an image preprocessing strategy that combines Task-driven Multi-granularity Resolution Adjustment (TMRA) with Prompt-guided Semantic-aware Cropping (PSC). This strategy allows the model to automatically perceive the task type and apply the appropriate processing method, thereby enhancing the accuracy of small object recognition and pixel-level segmentation, while simultaneously reducing computational redundancy by downsampling non-essential regions. Experimental results demonstrate that GeoMag not only achieves effective pixel-level processing but also delivers competitive performance across tasks of other granularities. These contributions provide a novel perspective for advancing remote sensing multimodal understanding toward a unified, all-granularity paradigm.

### v2

@@ Abstract.
The application of Vision-Language Models (VLMs) in remote sensing (RS) image understanding has achieved notable progress, demonstrating the basic ability to recognize and describe geographical entities. However, existing RS-VLMs are mostly limited to image-level and region-level tasks, lacking the capability to handle pixel-level tasks and performing poorly in small-object recognition scenarios. Moreover, RS-VLMs consume significant computational resources when processing high-resolution RS images, further restricting their practical applicability. In this context, we propose GeoMag (Geographical Magnifier), an end-to-end general-purpose large model framework for RS. GeoMag dynamically focuses the attention scope based on prompt semantics to effectively perform remote sensing image parsing across multiple levels of granularity. This method introduces Task-driven Multi-granularity Resolution Adjustment and Prompt-guided Semantic-aware Cropping, which adaptively reduce the spatial resolution of task-irrelevant regions while enhancing the visual representation of task-relevant areas. This approach improves the model’s perception of critical target regions, suppresses background redundancy, and reduces the computational cost of interpreting high-resolution RS imagery. Extensive comparative experiments on 10 benchmarks demonstrate that GeoMag not only excels in handling pixel-level tasks but also maintains competitive performance across tasks of other granularities compared to existing RS-VLMs.
VLMs; Remote Sensing Image; Pixel-level; Attention Mechanism.
copyright: acmlicensedjournalyear: 2025doi: 10.1145/XXXXXX.XXXXXXconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Ireland.booktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelandisbn: 979-8-4007-2035-2/2025/10ccs: Information systems Image searchccs: Information systems Multimedia and multimodal retrieval
@@ 1. Introduction
Figure 1. Overview of GeoMag’s capabilities. In addition to understanding both global and local image information for image-level tasks (green box) and region-level tasks (blue box), GeoMag extends the functionality of existing RS-VLMs by enabling pixel-level task processing—an ability typically absent in current models. As illustrated by the red box, GeoMag can automatically generate detailed descriptions of the image and produce segmentation masks for key geographical entities mentioned in the text. It can also perform user-directed segmentation based on specified prompts.
Remote sensing (RS) imagery, as a vital source of Earth observation data, serves scientific research, decision-making support, and the monitoring and evaluation of Sustainable Development Goals (SDGs) (Li et al., 2023). Therefore, effectively understanding the rich semantic content embedded in RS images becomes a fundamental step toward unlocking the value of RS data and promoting the intelligent development of earth science research (Ma et al., 2025, 2024).
@@ 2. Related Works
Pixel-level Parsing. In computer vision (Zeng et al., 2024; Yu et al., 2025; Hu et al., 2025), pixel-level RS image parsing tasks mainly include Referring Segmentation and Open Vocabulary Segmentation (Zhou et al., 2024; Pan et al., 2024; Liu et al., 2024c). To enable large language models (LLMs) to perform pixel-level tasks, PixelLM (Ren et al., 2024) and LISA (Lai et al., 2024) introduced a [SEG] token into the vocabulary of the LLM as a conditional input, allowing the decoder to generate segmentation masks from textual descriptions. GSVA (Xia et al., 2024) further introduced a [REJ] token to explicitly exclude irrelevant or unwanted targets, enabling the model to follow user intent more precisely. GLaMM (Rasheed et al., 2024) proposed a novel pixel-level task called Grounded Conversation Generation (GCG), aiming to identify and localize specific regions or objects in images through human-machine dialogue, thereby opening a promising new direction for pixel-level task research. However, these studies were conducted on natural images, which differ significantly from RS images in terms of semantic content and visual characteristics. Therefore, directly transferring such methods often fails to yield satisfactory results (Wang et al., 2023), and specialized optimization and adaptation for RS imagery are necessary.
@@ 3.2. MagCrop
To more effectively address challenges in remote sensing image parsing—such as varying granularity, sparse targets, and uneven spatial distributions—this work proposes an image preprocessing strategy named MagCrop. MagCrop reduces resource consumption (e.g., memory, GPU usage, and time) and enhances model response accuracy through the following principles: (1) For tasks emphasizing global features, it appropriately reduces image resolution. (2) For tasks with explicit spatial references (e.g., specific locations or objects), it preserves the resolution of task-relevant regions while compressing irrelevant areas. MagCrop integrates two key mechanisms: Task-driven Multi-granularity Resolution Adjustment and Prompt-guided Semantic-aware Cropping. These mechanisms help maintain the resolution of task-relevant regions while effectively reducing computational redundancy and emphasizing critical semantic content.
@@ 3.2.1. Task-driven Multi-granularity Resolution Adjustment
To address these issues, this study proposes a Task-driven Multi-granularity Resolution Adjustment strategy. This strategy dynamically adjusts the input image resolution based on the characteristics of each task, ensuring a balance between accuracy and computational efficiency.
@@
By adaptively adjusting the image input, Task-driven Multi-granularity Resolution Adjustment significantly improves GeoMag’s efficiency across tasks of different granularities, reduces the demand for computational resources.
@@ 3.2.2. Prompt-guided Semantic-aware Cropping
Current VLMs exhibit limited capability in perceiving and accurately parsing small visual details (Zhang et al., 2025; Kuckreja et al., 2024). To address the challenges RS-VLMs face in detecting small objects and to enhance their ability to focus on specific regions, GeoMag introduces Prompt-guided Semantic-aware Cropping. This strategy employs a gradient-weighted attention map generation method to query key regions in the image, emphasizing the deep semantic correlation between the prompt and the image content.
@@
By combining gradient weighting with the attention map generation strategy, Prompt-guided Semantic-aware Cropping enhances RS-VLM’s ability to focus on key regions and small objects within the image.
@@ 5. Conclusion
This study proposes GeoMag, a RS-VLM capable of handling multi-granularity RS image parsing tasks. By introducing components such as a Pixel Decoder, GeoMag extends its capabilities to pixel-level processing. It further presents MagCrop, an image preprocessing strategy that combines Task-driven Multi-granularity Resolution Adjustment with Prompt-guided Semantic-aware Cropping. This strategy allows the model to automatically perceive the task type and apply the appropriate processing method, thereby enhancing the accuracy of small object recognition and pixel-level segmentation, while simultaneously reducing computational redundancy by downsampling non-essential regions. Experimental results demonstrate that GeoMag not only achieves effective pixel-level processing but also delivers competitive performance across tasks of other granularities. These contributions provide a novel perspective for advancing remote sensing multimodal understanding toward a unified, all-granularity paradigm.